# On Bureau’s classification of quadratic differential equations in two variables free of movable critical points

### v2

@@ 3.7.3. Case $B\neq 0$, $n=-3$
and we thus find the necessary condition $D=0$ (Bureau’s analysis of this case fails to see that there are cases which are not free of movable critical points). We have
@@ 3.7.4. Case $B=0$
After solving for $y$, the system reduces to a linear inhomogeneous equation in $z$ with meromorphic coefficients. Its solutions may have movable critical points, arising from the poles of $y$. Its non-constant solutions are $y(t)=1/(t-t_{0})$,
@@
[…]
System III corresponds to the case $n=-1$. In it, by using a transformation $\text{T}_{1}$ instead of one of type $\text{T}_{2}$, we have obtained a normal form simpler than Bureau’s. The one appearing in Table 1 is obtained by changing the sign of $y$.
@@ 3.8. System XI
(This normal form is simpler than Bureau’s, and exhibits the natural symmetries of the system.) For $u=-1/y$, we have the canonical equation of index $1$
@@
from which we obtain the necessary condition $b=0$. By the symmetry of the system (27) consisting in exchanging $y$ and $z$, $a=0$ is a necessary condition as well. With these conditions,
$y^{\prime\prime}=2yy^{\prime}+qy^{\prime}+q^{\prime}y.$
The latter has the first integral $y^{\prime}=y^{2}+qy+K$, and thus $y$ solves a Riccati equation with holomorphic coefficients; the system is free of movable critical points.
@@ 3.9. System XII
$\begin{split}y^{\prime}&=y(2z+y)+2fy-a,\\ z^{\prime}&=-z(2y+z)-2fz+b.\end{split}$ (28)
For $u=-1/y$, we have
$\begin{split}u^{\prime}&=1-2fu-2uz-au^{2},\\ uz^{\prime}&=2z+bu-2fuz-uz^{2},\end{split}$
a canonical equation of index $2$, from which a necessary condition for the absence of movable critical points is found to be $b^{\prime}=0$. By the symmetries of (28), more evident when $z$ is replaced by $-z$, we also have the condition $a^{\prime}=0$. The involution $(y,z)\mapsto(-y,z+y+2f)$ preserves the general form of the systems (28), and acts upon their coefficients by
$(f,a,b)\mapsto(-f,-a,2f^{\prime}-a+b);$
in this way, we find that $f^{\prime\prime}=0$ is a third necessary condition for the absence of movable critical points. With these three, $y$ is a solution to Eq. (2) for the same $f$, for $\alpha=\frac{1}{2}a-b-f^{\prime}$ and $\beta=a$. Equation (2) is free of movable critical points: it is the general form of the equation leading to the Painlevé IV equation [Gam10, Ch. 1, §9]. When $f^{\prime}\neq 0$, up to translations in the independent variable and scalings, we may suppose that $f(t)=t$, recovering the Painlevé IV equation in its full generality. The cases where $f^{\prime}=0$ are integrated by elliptic functions: if $f$ is the constant $f_{0}$, Eq. (2) has the first integral $K$ defined by
$(y^{\prime})^{2}=y^{4}+4f_{0}y^{3}+4(f_{0}^{2}-\alpha)y^{2}+Ky+\beta^{2}$
(see also [Inc44, §14.331, case 6, p. 339]). In this case, (28) has the first integral
$y^{2}z+yz^{2}+2f_{0}yz-by-az-af_{0},$
whose generic fibers are elliptic fibrations over the domain of the independent variable.
System (28), with the integrability conditions, is birationally equivalent to the first-order rational system in two variables associated to Eq. (2), for, on the one hand, $y^{\prime}$ is a polynomial in $y$ and $z$, and, on the other, through (28), $z$ can be rationally expressed in terms of $y$ and $y^{\prime}$.
@@ 3.10. System XIII
(It is different from Bureau’s.) For $u=1/z$, we have the canonical equation of index $1$
@@
Hence, $a=0$ is a necessary condition for the absence of movable critical points. With it, $u=\frac{1}{2}y-p$ solves Eq. (22) for
@@
Following the discussion in Section 3.7.2.2, the necessary and sufficient conditions for Eq. (22) to be free of movable critical points are $f^{\prime\prime}=0$ and $g^{\prime}=0$, in which case we obtain the Painlevé II equation and some of its degenerations. Since, from (30),
$\begin{split}b&=f-2p^{\prime}+2p^{2},\\ p^{\prime\prime}&=2p^{3}+fp-g,\end{split}$
we have that, with the above conditions, $p$ is a solution to another instance of Eq. (22), one with the same $f$ and an opposite sign for $g$.
System (29), with the above integrability conditions, is birationally equivalent to the first-order system in two variables associated to Eq. (22), for $u$ and $u^{\prime}$ can be rationally expressed in terms of $y$ and $z$, and vice versa.
The conditions $f^{\prime\prime}=0$ and $g^{\prime}=0$ can be obtained through some canonical equations obtained from system (29) with $a=0$. For $u=2/y$, we have the canonical equation with index $3$
@@
from which we obtain the necessary condition for the absence of movable critical points $f^{\prime\prime}+2g^{\prime}=0$. The systems (29) with $a=0$ are preserved by the involution $(y,z)\mapsto(-y,z-y+4p)$, which acts upon their coefficients by $(p,b)\mapsto(-p,b+4p^{\prime})$. The involution acts upon $f$ and $g$ by $(f,g)\mapsto(f,-g)$. Hence, we also have the condition $f^{\prime\prime}-2g^{\prime}=0$. Together, these conditions give the previously discussed ones $f^{\prime\prime}=0$ and $g^{\prime}=0$.
@@ 3.11. System XIV
(different from Bureau’s). For $(u,v)=(1/y,z/y)$, we have the canonical equation of index $2$
@@
from which we obtain the necessary condition $b=0$. With it, for $u=1/z$, we have the canonical system of index $2$
@@
and, from it, a second necessary condition, $a^{\prime}=pa$, follows. Assuming these two, we find that $u=z+p$ is a solution to (23) for $q=\frac{1}{12}(p^{\prime}+p^{2}-a)$. By eliminating $a$ from $q$ and $q^{\prime}$, we have that $p$ is a particular solution to (23). System (31) and the first-order system associated to (23) are birationally equivalent. As discussed in Section 3.7.2.3, the necessary and sufficient condition for Eq. (23) to be free of movable critical points is $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$. Hence, for system (31), assuming the first two necessary conditions, this is a necessary and sufficient condition as well. System V, the one giving the first Painlevé equation, and system XIV are birationally equivalent.

### v3

@@ 3.7.3. Case $B\neq 0$, n=3n=-3
and we thus find the necessary condition $D=0$ (Bureau’s analysis of this case fails to see that there are instances in which the system is not free of movable critical points). We have
@@ 3.7.4. Case $B=0$
After solving for $y$, the system reduces to a linear inhomogeneous equation in $z$ with meromorphic coefficients. Its solutions may have movable critical points, arising from the poles of $y$. Its non-constant solutions are given by $y(t)=1/(t-t_{0})$ and
@@
Thus, for $n<0$, the system is always free of movable critical points. For $n>1$, the condition for the absence of movable critical points is $\mathrm{d}^{n-1}C/\mathrm{d}t^{n-1}=0$. This gives the normal forms for systems $\text{IX.A}_{0}(n)$ and $\text{IX.B}_{0}(n)$ in Table 1.
System III corresponds to the case n=1n=-1. In it, by using a transformation $\text{T}_{1}$ instead of one of type $\text{T}_{2}$, we have obtained a normal form simpler than Bureau’s. The one appearing in Table 1 is obtained by changing the sign of $y$.
@@ 3.8. System XI
(This normal form is simpler than Bureau’s, and exhibits the natural symmetries of the system.) For u=1/yu=-1/y, we have the canonical equation of index 11
@@
from which we obtain the necessary condition $b=0$. By the symmetry of the family (31) consisting in exchanging $y$ and $z$, we have that $a=0$ is a necessary condition as well. With these conditions, the system has the first integral $K=yz$, and reduces to the family of Riccati equations with holomorphic coefficients $y^{\prime}=y^{2}+qy-K$. It is free of movable critical points.
@@ 3.9. System XII
$\begin{split}y^{\prime}&=y(y-2z)-2fy+b,\\ z^{\prime}&=z(z-2y)+2fz-a.\end{split}$ (32)
This form is slightly different from Bureau’s, already at the level of the quadratic homogeneous part, which differs from the one in [Bur81b, §22] by a change of sign in $z$. System (32) is Hamiltonian: it may be written as $y^{\prime}=\partial\Phi/\partial z$, $z^{\prime}=-\partial\Phi/\partial y$, for
$\Phi=y^{2}z-yz^{2}-2fyz+ay+bz.$ (33)
For u=1/yu=-1/y, we have
$\begin{split}u^{\prime}&=1+2fu+bu^{2}+2uz,\\ uz^{\prime}&=2z-au+2fuz+uz^{2},\end{split}$
a canonical equation of index $2$, from which we obtain the necessary condition for the absence of movable critical points $a^{\prime}=0$. The symmetry of the family (32) consisting of interchanging $y$ and $z$ and acts upon its coefficients by
$(f,a,b)\mapsto(-f,-b,-a);$
from this, we obtain $b=0$ as a second necessary condition for the absence of movable critical points. Under the change of coordinates $(y,z)\mapsto(-y,z-y+2f)$, system (32) is mapped into one of the same family, with the transformation upon the coefficients given by
$(f,a,b)\mapsto(-f,a+b-2f^{\prime},-b).$
This is involutive. Through this, we find that $f^{\prime\prime}=0$ is a third necessary condition for the absence of movable critical points. With these three conditions, $z$ is a solution to Eq. (5) for $\alpha=\frac{1}{2}a+b-f^{\prime}$, $\beta=-\frac{1}{2}a^{2}$.
Eq. (5) is free of movable critical points: it is the general form of the equation leading to the Painlevé IV equation [Gam10, Ch. 1, §9]. When $f^{\prime}\neq 0$, up to translations in the independent variable and scalings, we may suppose that $f(t)=t$, recovering the Painlevé IV equation in its full generality. The cases where $f^{\prime}=0$ are integrated by elliptic functions: if $f$ is the constant $f_{0}$, Eq. (5) has the first integral $K$ defined by
$(y^{\prime})^{2}=y^{4}+4f_{0}y^{3}+4(f_{0}^{2}-\alpha)y^{2}+Ky-2\beta$
(see also [Inc44, §14.331, case 6, p. 339]). In this case, in which $f$ is constant, the system (32) is an autonomous one, and has in (33) a first integral; its generic level curves are elliptic.
System (32), with the integrability conditions, is birationally equivalent to the first-order system in two variables associated to the Painlevé IV equation (5); see Proposition 2. The Hamiltonian (33) is the one associated by Malmquist to the fourth Painlevé equation [Mal23, §33].
The above involutions generate a group of order six, isomorphic to the group $S_{3}$ of permutations in three symbols.
@@ 3.10. System XIII
(It is different from Bureau’s.) For $u=1/z$, we have the canonical equation of index 11
@@
Hence, $a=0$ is a necessary condition for the absence of movable critical points. With it, $\frac{1}{2}y-p$ solves Eq. (26) for
@@
Following the discussion in Section 3.7.2, the necessary and sufficient conditions for Eq. (26) to be free of movable critical points are $f^{\prime\prime}=0$ and $g^{\prime}=0$, in which case we obtain the Painlevé II equation and some of its degenerations. From (35),
$\begin{split}b&=f-2p^{\prime}+2p^{2},\\ p^{\prime\prime}&=2p^{3}+fp-g.\end{split}$
In particular, with the above conditions, $p$ is a solution to another instance of Eq. (26), one with the same $f$ but with an opposite sign for $g$.
The system is birationally equivalent to the first-order one associated to the Painlevé II equation (see Proposition 2).
The conditions $f^{\prime\prime}=0$ and $g^{\prime}=0$ can be obtained through some canonical equations obtained from system (34) with $a=0$. For $u=2/y$, we have the canonical equation with index 33
@@
from which we obtain the necessary condition for the absence of movable critical points $f^{\prime\prime}+2g^{\prime}=0$.
The systems (34) with $a=0$ are preserved by the involution $(y,z)\mapsto(-y,z-y+4p)$, which acts upon their coefficients by $(p,b)\mapsto(-p,b+4p^{\prime})$. The involution acts upon $f$ and $g$ by $(f,g)\mapsto(f,-g)$. Hence, we also have the necessary condition for the absence of movable critical points $f^{\prime\prime}-2g^{\prime}=0$. Together, these conditions give the previously discussed ones $f^{\prime\prime}=0$ and $g^{\prime}=0$.
@@ 3.11. System XIV
(different from Bureau’s). For $(u,v)=(1/y,z/y)$, we have the canonical equation of index 22
@@
from which we obtain the necessary condition $b=0$. With it, for $u=1/z$, we have the canonical system of index 22
@@
and, from it, a second necessary condition, $a^{\prime}=pa$, follows. Assuming these two, we find that $z+p$ is a solution to (27) for $q=\frac{1}{12}(p^{\prime}+p^{2}-a)$. By eliminating $a$ from $q$ and $q^{\prime}$, we have that $p$ is a particular solution to (27). System (36) and the first-order system associated to (27) are birationally equivalent. As discussed in Section 3.7.2, the necessary and sufficient condition for Eq. (27) to be free of movable critical points is $(q^{\prime\prime}-6q^{2})^{\prime\prime}=0$. Hence, for system (36), assuming the first two necessary conditions, this is a necessary and sufficient condition as well.
The system is birationally equivalent to the first-order one associated to the first Painlevé equation (cf. Proposition 2).
@@
Acknowledgments
The author thanks the referees for their helpful comments.